OpenStudy (mikeymaximum):

What is the prime factorization of 450? A. 2 • 3 • 5 B. 22 • 32 • 5 C. 2 • 32 • 52 D. 22 • 32 • 52

OpenStudy (matt101):

None of the answers you provided are right. B, C, and D all contain numbers that are not prime, and A doesn't multiply to give 450. Are the second 2's in the answer options meant to be exponents? If yes we can work with it! To find the prime factors of a number, divide the number into smaller and smaller parts until you only end up with the prime factors. For 450, it's easy to divide by 10: 45 * 10 = 450. 10 can be further factored as 5 * 2, which is the furthest you can go there since both those numbers are prime. 45 is 9 * 5, and 9 is 3 * 3, so 45 is really just 3 * 3 * 5. Putting it all together gives you: 450 = 10 * 45 = 2 * 5 * 3 * 3 * 5, making C your answer!

OpenStudy (fabulous18):

Answer: A